OpenStudy (anonymous):

Matthew currently has enough money to buy 45 toy cars. If the cost of each car was 10 cents less, Matthew could buy 5 more cars. How much money does Matthew have to spend on toy cars?

OpenStudy (anonymous):

CAN SOME SHOW ME A SIMPLE WAY TO THIS MATH PROBLEM

OpenStudy (lgbasallote):

let n = cost of toy car x = amt. matthew has 45n < x 45+5(n-10) < x make sense? hehe

OpenStudy (anonymous):

Matt's money in the first case is 45n, where n is the cost of the toy car. In the second case, its 50(n - 0.10). 50(n - 0.10) = 45n 50n - 5 = 45n n = 1

OpenStudy (anonymous):

So Matt has 45 dollars to begin with

OpenStudy (anonymous):

Let t be the original cost of a toy car in cents or pennys. 45 t is the total amount of money on hand to buy toy cars. Solve the following equation for t, the original cost per car.\[45 t=50(t-10) \]\[t=100 \]\[45t = 45*100 = 4500 \text{ cents} \text{ or } \text{$\$$45} \]